# TriG-NER: Triplet-Grid Framework for Discontinuous Named Entity Recognition

### v2

@@ Abstract.
Discontinuous Named Entity Recognition (DNER) presents a challenging problem where entities may be scattered across multiple non-adjacent tokens, making traditional sequence labelling approaches inadequate. Existing methods predominantly rely on custom tagging schemes to handle these discontinuous entities, resulting in models tightly coupled to specific tagging strategies and lacking generalisability across diverse datasets. To address these challenges, we propose TriG-NER, a novel Triplet-Grid Framework that introduces a generalisable approach to learning robust token-level representations for discontinuous entity extraction. Our framework applies triplet loss at the token level, where similarity is defined by word pairs existing within the same entity, effectively pulling together similar and pushing apart dissimilar ones. This approach enhances entity boundary detection and reduces the dependency on specific tagging schemes by focusing on word-pair relationships within a flexible grid structure. We evaluate TriG-NER on three benchmark DNER datasets and demonstrate significant improvements over existing grid-based architectures. These results underscore our framework’s effectiveness in capturing complex entity structures and its adaptability to various tagging schemes, setting a new benchmark for discontinuous entity extraction.
Discontinuous Named Entity Recognition, Medical Named Entity Recognition, Medical Text Mining
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Information extractionccs: Applied computing Health informatics
@@ 1. Introduction
Figure 1. A Case example involving discontinuous mentions with Gold Standard (a) Our proposed TriG-NER enables to perfectly extract the DNE (b,c,d) LLMs face challenges in DNER as those are primarily trained to capture continuous sequences of text, making it difficult for them to recognise entities split across discontinuous regions while maintaining coherence in prediction.
@@ 2. Related Works
2.1. Discontinuous Named Entity Recognition
Named entity extraction and recognition has traditionally been viewed as a sequence labelling task using the Begin-Inside-Outside (BIO) tags; however, this traditional approach fails for more complex entities such as discontinuous entities. Researchers have recently focused on improving discriminative discontinuous entity recognition through various tagging schemes and methods. Tang et al. (2015) (Tang et al., 2015) was the first to extend BIO sequential tagging to BIOHD to distinguish inter-entity boundaries, which subsequent studies (Metke-Jimenez and Karimi, 2016; Tang et al., 2018) followed. More recently, Corro (2024) (Corro, 2024) proposed a two-layer tagging scheme that uses ten tags; however, these methods fail to capture complex discontinuous entities and suffer from decoding ambiguity. Span-based methods (Wang and Lu, 2019; Li et al., 2021; Huang et al., 2023) typically involve the identification of all candidate spans and the merging of disjoint spans. The two-step process, however, is vulnerable to error propagation and identifying all possible span candidates is resource-exhaustive. Other discriminative methods, such as hypergraphs (Muis and Lu, 2016; Wang and Lu, 2018) and stack-and-buffer transitions (Dai et al., 2020), are also explored yet still suffer from error propagation. On the other hand, generative methods (Fei et al., 2021; Yan et al., 2021; Zhang et al., 2022; Xia et al., 2023), leverage sequence-to-sequence language models to directly generate entity spans and types that overcome the challenges presented by different complex entity structures. With the advent of ChatGPT, research in applying large language model (LLM) prompting to discontinuous NER has also seen increased attention (Zhu et al., 2024; Zhao et al., 2024). However, generative models are optimised for next-word prediction, not NER, predisposing it to incorrect biases.
@@ 3.2. Word-Pair Relationship Grid
The input sentence is first passed through an encoder layer, where we utilise pre-trained language models (PLMs) such as BERT (Devlin et al., 2019), BioClinicalBERT (Alsentzer et al., 2019), PharmBERT (ValizadehAslani et al., 2023), and PubMedBERT (Gu et al., 2021). These models generate contextualised word embeddings $\textbf{H}^{em}\in\mathbb{R}^{n\times d}$, where $d$ is the embedding dimension. A bidirectional LSTM layer is then applied to capture sequential dependencies in the sentence. The embeddings are then passed through two distinct modules: a Convolution Layer and a Biaffine transformation. The Convolution Layer generates enhanced word-pair representations $\textbf{H}^{co}\in\mathbb{R}^{n\times n\times d^{co}}$, where $d^{co}$ is the convolution dimension, while the Biaffine transformation computes word-pair relationships $\textbf{H}^{bi}\in\mathbb{R}^{n\times n\times d^{bi}}$. These representations are combined in a Co-Predictor Layer, where a linear layer and an MLP map $\textbf{H}^{bi}$ and $\textbf{H}^{co}$ to tag relation logits $\textbf{Y}^{bi}$ and $\textbf{Y}^{co}\in\mathbb{R}^{n\times n\times c}$. The final grid tag logits are obtained by combining the two: $\textbf{Y}=\textbf{Y}^{bi}+\textbf{Y}^{co}$.
@@ 3.3. Grid Tagging and Decoding
The grid tagging system classifies word-pair relationships using three tag classes: None, Next-Neighboring-Word (NNW), and Tail-Head-Word (THW). These classes define whether a word pair has no relationship, a neighbouring relationship within an entity, or represents the start and end of an entity, respectively. Once word-pair relationships are classified, the grid decoding process begins, which is crucial for discontinuous entity extraction. The system takes the final grid tag logits Y and decodes the predicted relationships into entity structures. By focusing on word pairs rather than individual tokens, the grid structure allows our model to flexibly identify discontinuous entity boundaries, which are common in complex entity recognition tasks. The grid tagging and decoding approach enables the model to handle non-contiguous entity spans by considering the relationships between word pairs, making it robust against the limitations of sequential tagging schemes.
@@ 3.4.2. Word-Pair Grid Implementation
We extract our triplets from the word-pair grid representations in our framework. Unlike most sample-based triplet loss implementations that define similarity by sample classes, we define the similarity of our triplet elements based on their existence within entities. For the anchor candidates, we use word-pair grid points that exist in any entity. Each anchor candidate is then matched with positive and negative candidates. Positive candidates are word pairs that co-exist with the anchor in any entities, while negative candidates are word pairs that don’t belong to any entity the anchor is a part of. We illustrate this candidate selection in Figure 3.
@@ 3.4.3. Triplet Selection
Due to the exponential increase of positive and negative candidates as the sample length increases, we further limit the positive and negative candidate selection by using a candidate window centred on the anchor and by specifically using unique anchor-positive pairs. The unique anchor-positive pair setup utilises only the top half triangle of the grid (Figure 3) where an anchor token pair $tp_{1}$ is paired with a positive candidate $tp_{2}$, but when $tp_{2}$ is set as an anchor, $tp_{1}$ will not be considered as a positive candidate anymore. This reduces possible redundant information that is not helpful for training while simultaneously reducing the number of triplets. A comparison of performance between unique and non-unique anchor-positive pairs is provided in Table 5.
@@ 4.1. Datasets
Following previous studies on discontinuous named entity recognition, we use three datasets in the biomedical domain to assess the performance of our proposed system. The CSIRO Adverse Drug Event Corpus (CADEC) (Karimi et al., 2015) is a collection of medication consumer posts annotated for entity identification from the public forum AskAPatient. We follow previous literature and use only the adverse drug reaction (ADR) entities. ShARe13 (Pradhan et al., 2013) and ShARe14 (Mowery et al., 2014) datasets are part of the Shared Annotated Resources used for the CLEF eHealth Challenge in 2013 and 2014, respectively. They consist of clinical reports annotated for the identification and normalisation of disease disorders. For all datasets, we use the sentence-based preprocessing script and dataset splits provided by Dai et al. (Dai et al., 2020) and convert the produced inline format to JSON [^1] Script provided in the code repository. following Li et al. (Li et al., 2022). Table 2 shows each dataset’s statistics.
Table 2. Data statistics
CADEC ShARe13 ShARe14
Total Sentences 7,597 18,767 34,618
Total Entities 6,318 11,148 19,073
Continuous Entities 5,639 10,060 17,417
- Percentage 89.25% 90.24% 91.32%
- Number of tokens 1-36 1-9 1-9
Disc. Entities 679 1,088 1,658
- Percentage 10.75% 9.76% 8.68%
- Number of tokens 2-13 2-7 2-7
- Start-End Distance 3-20 3-23 3-23
@@ 4.3. Implementation Details
We evaluate our framework using the established training, validation, and test splits by (Dai et al., 2020). We list best-performing model setups for each dataset in the Appendix D. Each model is trained using the AdamW optimiser with a learning rate of 5e-4 for a maximum of 60 epochs and an early stop of 10 epochs. We take the best-performing model on the validation set based on the micro-F1 score. We use a batch size of 12, 6, and 6 for CADEC, ShARe13, and ShARe14, respectively. Our best setup for the CADEC dataset uses a fine-tuned BioBERT, while both ShARe datasets achieve better results with fine-tuned PubMedBERT. A comparison of PLMs is provided in Table 7. All models are trained using an NVIDIA RTX A4500.
@@ 5.1. Overall Performance
A comprehensive evaluation of our framework compared to other studies is provided in Table 3. The results reflect the performance of our framework on the entire test set, as well as on discontinuous elements, with isolated evaluations on sentences containing at least one discontinuous entity (DiscSent) and on discontinuous entities exclusively (DiscEnt). Our framework demonstrates a clear improvement in both F1 score and precision over W2NER, the best-performing baseline method. The ShARe14 dataset shows the most significant improvement in F1 score, with a 1.23% increase, reaching 82.54. Similarly, the CADEC and ShARe13 datasets show increases of 0.76% (73.43) and 1.06% (83.22), respectively. Furthermore, our framework outperforms the baseline models when focusing on discontinuous elements, with improvements of 0.79%, 0.63%, and 3.19% for DiscSent, and 3.98%, 2.68%, and 5.13% for DiscEnt across the CADEC, ShARe13, and ShARe14 datasets, respectively. Complete performance metrics may be found in Appendix A. These results underscore the strength of our TriG-NER framework in capturing the complexities of discontinuous entities by leveraging word-pair similarities and dissimilarities. By focusing on token-level relationships within a flexible grid structure, our approach demonstrates superior performance in both overall entity recognition and specifically in handling discontinuous elements, highlighting its adaptability and effectiveness compared to traditional methods.
@@ 5.3. Window Size
5.4. Encoder Language Models
We evaluated the performance of our framework with different pre-trained language models for the encoder, using the best-performing model setup for each dataset. Table 7 presents the results for four biomedical BERT variants, both with and without our grid-based triplet framework. Overall, BioBERT yields the best results for the CADEC dataset, while PubMedBERT outperforms others for both ShARe datasets. The application of our framework further enhances these scores by 0.93%, 1.22%, and 1.12%, respectively, demonstrating that our framework effectively captures local dependencies via the word-pair triplet implementation. Additionally, our framework consistently improves the performance of most PLMs tested, with the exception of BioClinicalBERT for CADEC and ShARe14. In Table 8, we present the performance improvements achieved by finetuning the pre-trained language models using a next-word prediction task for each dataset. As expected, finetuning enhances the scores across the board, with more pronounced improvements observed in the ShARe datasets, likely due to the specialised clinical terminology in those datasets compared to the more natural language used in online forums like CADEC.
@@ 5.4. Encoder Language Models
5.5. Qualitative Analysis
In this section, we demonstrate the effectiveness of our word-pair grid-based triplet framework through a qualitative analysis of the extracted entities, comparing the results with those of the best-performing baseline model and LLMs, such as Gemini 1.5-flash (Team et al., 2023) and GPT-4o (Achiam et al., 2023). We trained and fine-tuned both our model and the replicated baseline model using tokenised sentences as direct inputs, while the LLMs were not fine-tuned and were provided with task-specific prompts that described the task, input, and expected output format. For few-shot prompts, we included two examples from the training data. Table 10 presents a case study based on a CADEC sample, with additional case studies and prompt templates available in Appendix E and Appendix F.
While our framework uses the same tags as W2NER, it goes further by leveraging word-pair relationships to accurately recognise multiple non-adjacent entity segments within the input text. In contrast, W2NER processes word pairs in isolation, which limits its ability to recognise entities with more than two disjoint spans, such as ”Pain in my lower legs” and ”cramping in my lower legs”, indexed as [0, 3, 4, 7, 8] and [2, 3, 4, 7, 8], respectively. Furthermore, W2NER struggles to detect uncommon, domain-specific terms and abbreviations, particularly when the entity consists of just one word. For example, in Figure F4, our framework successfully extracts the entity ”PFO”, which stands for ”Patent Foramen Ovale”, despite the presence of other domain-specific terms. By contrast, W2NER incorrectly extracts ”MV”, which in this context likely refers to ”mitral valve”, but is not a disorder.
With LLMs’ recent success and popularity for general language generation tasks, we evaluate their performance in extracting entity indexes through zero-shot and few-shot chain-of-thought (CoT) prompting. Because LLMs are optimised for next-word prediction, these models are prone to alignment and indexing problems where, despite clear instructions, the indexes returned do not correspond to the entity words identified. We found that explicitly including the entity words in the return format prompt helps partially but does not entirely resolve the problem. For instance, in Figure F2, the entity words ”loss of range of motion” are correctly identified; however, the indexes provided are one or two positions off. In some cases, the number of words identified does not equate to the number of indexes returned, such as ”{‘entity’: ‘loss of range of motion’, ‘index’: [32, 36], ‘type’: ‘ADR’}”. Furthermore, LLMs fail to extract discontinuous entities most of the time. In Table 10, both Gemini and GPT-4o completely missed the overlapping continuous and discontinuous entities in the sample despite identifying relevant parts such as ”Pain and cramping”, ”hands”, and ”lower legs”. They cannot effectively split and combine disjoint spans to form discontinuous entities such as ”Pain in my hands” and ”Pain in my lower legs”. GPT-4o Few-shot CoT goes as far as returning the whole input instead of associating the relevant spans together. Lastly, LLMs are prone to extracting entities unrelated to the entity type provided. For instance, body parts such as ”hands” and ”lower legs” in Table 10 and medical procedures such as ”CABG” (coronary artery bypass graft surgery) are separately identified as ADRs and Disorders, respectively.
@@ 5.5. Qualitative Analysis
Table 10. Case study for CADEC comparing the entity extraction results from trained models using our TriG-NER framework, a baseline model, and LLMs employing zero-shot and few-shot chain-of-thought (CoT) prompt engineering. The table compares how each method identifies discontinuous entities within a sample sentence from the CADEC. The models trained with our framework demonstrate more accurate entity recognition, especially for non-adjacent entity segments. Prompt templates for the LLMs are provided in Appendix E. Green highlight indicates correctly identified entities. Red highlight indicates otherwise.

### v3

@@ Abstract.
Discontinuous Named Entity Recognition (DNER) presents a challenging problem where entities may be scattered across multiple non-adjacent tokens, making traditional sequence labelling approaches inadequate. Existing methods predominantly rely on custom tagging schemes to handle these discontinuous entities, resulting in models tightly coupled to specific tagging strategies and lacking generalisability across diverse datasets. To address these challenges, we propose TriG-NER, a novel Triplet-Grid Framework that introduces a generalisable approach to learning robust token-level representations for discontinuous entity extraction. Our framework applies triplet loss at the token level, where similarity is defined by word pairs existing within the same entity, effectively pulling together similar and pushing apart dissimilar ones. This approach enhances entity boundary detection and reduces the dependency on specific tagging schemes by focusing on word-pair relationships within a flexible grid structure. We evaluate TriG-NER on three benchmark DNER datasets and demonstrate significant improvements over existing grid-based architectures. These results underscore our framework’s effectiveness in capturing complex entity structures and its adaptability to various tagging schemes, setting a new benchmark for discontinuous entity extraction. [^1] Code available at https://github.com/adlnlp/trig_ner.
Discontinuous Named Entity Recognition; Medical Named Entity Recognition; Medical Text Mining
copyright: acmlicensedjournalyear: 2025copyright: ccconference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714639isbn: 979-8-4007-1274-6/25/04ccs: Information systems Information extractionccs: Applied computing Health informatics
Figure 1. A case example involving discontinuous mentions. TriG-NER (a) perfectly extracts the DNE. LLMs (b,c,d), primarily trained to capture continuous sequences of text, face challenges in recognising entities split across discontinuous regions while maintaining coherence in prediction.
@@ 2. Related Works
2.1. Discontinuous Named Entity Recognition
Named entity extraction and recognition has traditionally been viewed as a sequence labelling task using the Begin-Inside-Outside (BIO) tags; however, this traditional approach fails for more complex entities such as discontinuous entities. Researchers have recently focused on improving discriminative discontinuous entity recognition through various tagging schemes and methods. Tang et al. (2015) (Tang et al., 2015) was the first to extend BIO sequential tagging to BIOHD to distinguish inter-entity boundaries, which subsequent studies (Metke-Jimenez and Karimi, 2016; Tang et al., 2018) followed. More recently, Corro (2024) (Corro, 2024) proposed a two-layer tagging scheme that uses ten tags; however, these methods fail to capture complex discontinuous entities and suffer from decoding ambiguity. Span-based methods (Wang and Lu, 2019; Li et al., 2021; Huang et al., 2023) involve the identification of all candidate spans and the merging of disjoint spans. The two-step process, however, is vulnerable to error propagation and identifying all possible span candidates is resource-exhaustive. Other discriminative methods, such as hypergraphs (Muis and Lu, 2016; Wang and Lu, 2018) and stack-and-buffer transitions (Dai et al., 2020), are also explored yet still suffer from error propagation. On the other hand, generative methods (Fei et al., 2021; Yan et al., 2021; Zhang et al., 2022; Xia et al., 2023), leverage sequence-to-sequence language models to directly generate entity spans and types that overcome the challenges presented by different complex entity structures. With the advent of ChatGPT, research in applying large language model (LLM) prompting to discontinuous NER has also seen increased attention (Zhu et al., 2024; Zhao et al., 2024). However, generative models are optimised for next-word prediction, not NER, predisposing it to incorrect biases.
@@ 3.2. Word-Pair Relationship Grid
The input sentence is first passed through an encoder layer, where we utilise pre-trained language models (PLMs) such as BERT (Devlin et al., 2019), BioClinicalBERT (Alsentzer et al., 2019), PharmBERT (ValizadehAslani et al., 2023), and PubMedBERT (Gu et al., 2021). These models generate contextualised word embeddings $\textbf{H}^{em}\in\mathbb{R}^{n\times d}$, where $d$ is the embedding dimension. A bidirectional LSTM layer is then applied to capture sequential dependencies in the sentence. The embeddings are then passed through two distinct modules: a Convolution Layer and a Biaffine transformation. The Convolution Layer module (Li et al., 2022) generates enhanced word-pair representations $\textbf{H}^{co}\in\mathbb{R}^{n\times n\times d^{co}}$, where $d^{co}$ is the convolution dimension, while the Biaffine transformation computes word-pair relationships $\textbf{H}^{bi}\in\mathbb{R}^{n\times n\times d^{bi}}$. These representations are combined in a Co-Predictor Layer, where a linear layer and an MLP map $\textbf{H}^{bi}$ and $\textbf{H}^{co}$ to tag relation logits $\textbf{Y}^{bi}$ and $\textbf{Y}^{co}\in\mathbb{R}^{n\times n\times c}$. The final grid tag logits are obtained by combining the two: $\textbf{Y}=\textbf{Y}^{bi}+\textbf{Y}^{co}$.
@@ 3.3. Grid Tagging and Decoding
The grid tagging system, reproduced from (Li et al., 2022), classifies word-pair relationships using three tag classes: None, Next-Neighboring-Word (NNW), and Tail-Head-Word (THW). These classes define whether a word pair has no relationship, a neighbouring relationship within an entity, or represents the start and end of an entity, respectively. Once word-pair relationships are classified, the grid decoding process begins, which is crucial for discontinuous entity extraction. The system takes the final grid tag logits Y and decodes the predicted relationships into entity structures. By focusing on word pairs rather than individual tokens, the grid structure allows our model to flexibly identify discontinuous entity boundaries, which are common in complex entity recognition tasks. The grid tagging and decoding approach enables the model to handle non-contiguous entity spans by considering the relationships between word pairs, making it robust against the limitations of sequential tagging schemes.
@@ 3.4.2. Word-Pair Grid Implementation
We extract our triplets from the word-pair grid representations in our framework. Unlike most sample-based triplet loss implementations that define similarity by sample classes, we define the similarity of our triplet elements based on their existence within entities. For the anchor candidates, we use word-pair grid points that exist in any entity. Each anchor candidate is then matched with positive and negative candidates. Positive candidates are word pairs that co-exist with the anchor in any entity, while negative candidates are word pairs that don’t belong to any entity the anchor is a part of. We illustrate this candidate selection in Figure 3.
@@ 3.4.3. Triplet Selection
Due to the exponential increase of positive and negative candidates as the sample length increases, we further limit the positive and negative candidate selection by using a candidate window centred on the anchor and by specifically using unique anchor-positive pairs. The unique anchor-positive pair setup utilises only the top half triangle of the grid (Figure 3) where an anchor token pair $tp_{1}$ is paired with a positive candidate $tp_{2}$, but when $tp_{2}$ is set as an anchor, $tp_{1}$ will not be considered as a positive candidate anymore. This reduces possible redundant information that is not helpful for training, while simultaneously reducing the number of triplets. A comparison of performance between unique and non-unique anchor-positive pairs is provided in Table 4.
@@ 4.1. Datasets
Following previous studies on discontinuous named entity recognition, we use three datasets in the biomedical domain to assess the performance of our proposed system. The CSIRO Adverse Drug Event Corpus (CADEC) [^2] https://doi.org/10.4225/08/570FB102BDAD2 (Karimi et al., 2015) is a collection of medication consumer posts annotated for entity identification from the public forum AskAPatient. We follow previous literature and use only the adverse drug reaction (ADR) entities. ShARe13333https://doi.org/10.13026/rxa7-q798 (Pradhan et al., 2013) and ShARe14444https://doi.org/10.13026/0zgk-9j94 (Mowery et al., 2014) datasets are part of the Shared Annotated Resources used for the CLEF eHealth Challenge in 2013 and 2014, respectively. They consist of clinical reports annotated for the identification and normalisation of disease disorders. For all datasets, we use the sentence-based, token-level preprocessing script and dataset splits provided by Dai et al. (Dai et al., 2020) and convert the produced inline format to JSON following Li et al. (Li et al., 2022). Detailed statistics can be found in Appendix A.
@@ 4.3. Implementation Details
We evaluate our framework using the established training, validation, and test splits by (Dai et al., 2020). We list best-performing model setups for each dataset in the Appendix E. Each model is trained using the AdamW optimiser with a learning rate of 5e-4 for a maximum of 60 epochs and an early stop of 10 epochs. We take the best-performing model on the validation set based on the micro-F1 score. We use a batch size of 12, 6, and 6 for CADEC, ShARe13, and ShARe14, respectively. Our best setup for the CADEC dataset uses a fine-tuned BioBERT, while both ShARe datasets achieve better results with fine-tuned PubMedBERT. A comparison of PLMs is provided in Table 6. All models are trained using an NVIDIA RTX A4500.
@@ 5.1. Overall Performance
A comprehensive evaluation of our framework compared to other studies is provided in Table 2. The results reflect the performance of our framework on the entire test set, as well as on discontinuous elements, with isolated evaluations on sentences containing at least one discontinuous entity (DiscSent) and on discontinuous entities exclusively (DiscEnt). Our framework demonstrates a clear improvement in both F1 score and precision over W2NER, the best-performing baseline method. The ShARe14 dataset shows the most significant improvement in F1 score, with a 1.23% increase, reaching 82.54. Similarly, the CADEC and ShARe13 datasets show increases of 0.76% (73.43) and 1.06% (83.22), respectively. Furthermore, our framework outperforms the baseline models when focusing on discontinuous elements, with improvements of 0.79%, 0.63%, and 3.19% for DiscSent, and 3.98%, 2.68%, and 5.13% for DiscEnt across the CADEC, ShARe13, and ShARe14 datasets, respectively. Complete performance metrics may be found in Appendix B. These results underscore the strength of our TriG-NER framework in capturing the complexities of discontinuous entities by leveraging word-pair similarities and dissimilarities. By focusing on token-level relationships within a flexible grid structure, our approach demonstrates superior performance in both overall entity recognition and specifically in handling discontinuous elements, highlighting its adaptability and effectiveness compared to traditional methods.
@@ 5.2. Triplet Selection
5.3. Window Size
@@
5.4. Encoder Language Models
We evaluated the performance of our framework with different pre-trained language models for the encoder, using the best-performing model setup for each dataset. Table 6 presents the results for four biomedical BERT variants, both with and without our grid-based triplet framework. Overall, BioBERT yields the best results for the CADEC dataset, while PubMedBERT outperforms others for both ShARe datasets. The application of our framework further enhances these scores by 0.93%, 1.22%, and 1.12%, respectively, demonstrating that our framework effectively captures local dependencies via the word-pair triplet implementation. Additionally, our framework consistently improves the performance of most PLMs tested, with the exception of BioClinicalBERT for CADEC and ShARe14. In Table 7, we present the performance improvements achieved by finetuning the PLMs on a next-word prediction task for each dataset. As expected, finetuning enhances the scores across the board, with more pronounced improvements observed in the ShARe datasets, likely due to the specialised clinical terminology in those datasets compared to the more natural language used in online forums like in CADEC.
@@ 5.4. Encoder Language Models
5.5. Qualitative Analysis
In this section, we demonstrate the effectiveness of our word-pair grid-based triplet framework through a qualitative analysis of the extracted entities, comparing the results with the best-performing baseline model and LLMs, such as Gemini 1.5-flash (Team et al., 2023) and GPT-4o (Achiam et al., 2023). We trained and fine-tuned both our model and the replicated baseline model using tokenised sentences as direct inputs, while the LLMs were not fine-tuned and were provided with task-specific prompts that described the task, input, and expected output format. For few-shot prompts, we included two examples from the training data. Table 9 presents a case study from CADEC. Additional case studies and prompt templates are available in Appendix F and G.
While our framework uses the same tags as W2NER (Li et al., 2022), it goes further by leveraging word-pair relationships to accurately recognise multiple non-adjacent entity segments within the input text. In contrast, W2NER processes word pairs in isolation, which limits its ability to recognise entities with more than two disjoint spans, such as ”Pain in my lower legs” and ”cramping in my lower legs”, indexed as [0, 3, 4, 7, 8] and [2, 3, 4, 7, 8], respectively. Furthermore, W2NER struggles to detect uncommon, domain-specific terms and abbreviations, particularly when the entity consists of just one word. For example, in Figure G4, our framework successfully extracts the entity ”PFO”, which stands for ”Patent Foramen Ovale”, despite the presence of other domain-specific terms. In contrast, W2NER incorrectly extracts ”MV”, which likely refers to ”mitral valve”, but is not a disorder.
With LLMs’ recent success and popularity for general language generation tasks, we evaluate their performance in extracting entity indexes through zero-shot and few-shot chain-of-thought (CoT) prompting. Because LLMs are optimised for next-word prediction, these models are prone to alignment and indexing problems where, despite clear instructions, the indexes returned do not correspond to the entity words identified. We found that explicitly including the entity words in the return format prompt helps partially but does not entirely resolve the problem. For instance, in Figure G2, the entity words ”loss of range of motion” are correctly identified; however, the indexes provided are one or two positions off. In some cases, the number of words identified does not equate to the number of indexes returned, such as ”{‘entity’: ‘loss of range of motion’, ‘index’: [32, 36], ‘type’: ‘ADR’}”. Furthermore, LLMs fail to extract discontinuous entities most of the time. In Table 9, both Gemini and GPT-4o completely missed the overlapping continuous and discontinuous entities in the sample despite identifying relevant parts such as ”Pain and cramping”, ”hands”, and ”lower legs”. They cannot effectively split and combine disjoint spans to form discontinuous entities such as ”Pain in my hands” and ”Pain in my lower legs”. GPT-4o Few-shot CoT goes as far as returning the whole input instead of associating the relevant spans together. Lastly, LLMs are prone to extracting entities unrelated to the entity type provided. For instance, body parts such as ”hands” and ”lower legs” in Table 9 and medical procedures such as ”CABG” (coronary artery bypass graft surgery) in Figure G4 are separately identified as ADRs and Disorders, respectively.
Table 9. Case study for CADEC comparing the entity extraction results from trained models using our TriG-NER framework, a baseline model, and LLMs employing zero-shot and few-shot chain-of-thought (CoT) prompt engineering. The table compares how each method identifies discontinuous entities within a sample sentence from the CADEC. The models trained with our framework demonstrate more accurate entity recognition, especially for non-adjacent entity segments. Prompt templates for the LLMs are provided in Appendix F. Green highlight indicates correctly identified entities. Red highlight indicates otherwise.